What's is the magnetic field at the center of the coil of radius 5×10^-2 m carrying a current of 0. 25 A and having 40 turns


1
Expert's answer
2021-04-26T17:13:38-0400
"B=\frac{\mu_0NI}{2R}\\B=\frac{(40)(0.25)(4\pi\cdot10^{-7})}{2(0.05)}=1.26\cdot10^{-4}\ T"